OpenStudy (anonymous):

What is the exact value of: tan^-1(tan(4pi/5)) I put 4pi/5 as the answer, why is that wrong?

OpenStudy (kropot72):

\[\tan (\frac{4 \pi}{5})=-0.726542528\] \[\tan^{-1} -0.726542528=-0.6283185307\ radians=-36\ degrees\]

OpenStudy (anonymous):

it is wrong because \(\frac{4\pi}{5}>\frac{\pi}{2}\)

OpenStudy (anonymous):

you want the same angle (number) that will give you the same tangent, only it must be between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\) because that is the range of arctangent

OpenStudy (anonymous):

that is why the obvious answer is not correct

OpenStudy (anonymous):